# Fractal counterparts of density combinatorics theorems in continued fractions

### v1

@@ Abstract.
We build a bridge from density combinatorics to dimension theory of continued fractions. We establish a fractal transference principle that transfers common properties of subsets of $\mathbb{N}$ with positive upper density to properties of subsets of irrationals in $(0,1)$ for which the set $\{a_{n}(x)\colon n\in\mathbb{N}\}$ of partial quotients induces an injection $n\in\mathbb{N}\mapsto a_{n}(x)\in\mathbb{N}$. Let ()(*)( ∗ ) be a certain property that holds for any subset of $\mathbb{N}$ with positive upper density. The principle asserts that for any subset $S$ of $\mathbb{N}$ with positive upper density, there exists a set $E_{S}$ of Hausdorff dimension $1/2$ such that the set $\bigcup_{n\in\mathbb{N}}\bigcap_{x\in E_{S}}\{a_{n}(x)\}\cap S$ has the same upper density as that of $S$, and thus inherits property ()(*)( ∗ ). Examples of ()(*)( ∗ ) include the existence of arithmetic progressions of arbitrary lengths and the existence of arbitrary polynomial progressions, known as Szemerédi’s and Bergelson-Leibman’s theorems respectively. In the same spirit, we establish a relativized version of the principle applicable to the primes, to the primes of the form $y^{2}+z^{2}+1$, to the sets given by the Piatetski-Shapiro sequences.
@@ 1. Introduction
$\overline{d}(S)=\limsup_{N\to\infty}\frac{1}{N}\#({S\cap[1,N]}),$
@@
$d_{\rm B}(S)=\limsup_{N\to\infty}\frac{1}{N}\sup_{M\in\mathbb{N}}\#({S\cap[M,M+N]}).$
Note that $d_{\rm B}(S)\geq\overline{d}(S)$. The positive upper density assumption in Szemerédi’s theorem can actually be weakened to the positivity of upper Banach density.
Furstenberg [10, 11] established the multiple recurrence theorem and the correspondence principle, linking the field of density combinatorics to recurrence in ergodic theory. Using this principle he obtained a new proof of Szemerédi’s theorem. This development laid a foundation for further extensions of Szemeredi’s theorem to several different directions, see [2, 12, 14, 34, 36] for example. The aim of this paper is to provide counterparts of these density combinatorics theorems in dimension theory of continued fractions.
Each $x\in(0,1)\setminus\mathbb{Q}$ has the unique infinite regular continued fraction expansion
[…]
@@
$E=\{x\in(0,1)\setminus\mathbb{Q}\colon a_{m}(x)\neq a_{n}(x)\ \text{\rm for all }m,n\in\mathbb{N},\ m\neq n\}.$
An investigation of this set was suggested by Erdős in the early stage of dimension theory of continued fractions. Ramharter [31] proved that $E$ is of Hausdorff dimension $1/2$. Ramharter’s result is a refinement of Good’s landmark theorem [13] which states that the set $\{x\in(0,1)\setminus\mathbb{Q}\colon a_{n}(x)\to\infty\text{ as }n\to\infty\}$ is of Hausdorff dimension $1/2$.
@@ 1.1. Statements of results and their consequences
We now state our two main results (Theorem 1.1 and Theorem 1.4), along with five consequential theorems which are manifestations of density combinatorics theorems in the regular continued fraction.
@@ 1.1.1. First main result
Equation (1.3) is a ‘pointwise version’ of the second equation in (1.2). Equation (1.3) asserts that the condition $\overline{d}(\{a_{n}(x)\colon n\in\mathbb{N}\}\cap S)=\overline{d}(S)$ does not cause a dimension drop in $E$. More importantly, it allows us to immediately transfer statements on subsets of $\mathbb{N}$ with positive upper density to statements on continued fractions of points in sets of Hausdorff dimension $1/2$. For example, from (1.3) and Szemerédi’s theorem it follows that if $S\subset\mathbb{N}$ has positive upper density then for any $x\in(0,1)\setminus\mathbb{Q}$ in a set of Hausdorff dimension $1/2$ and for every integer $\ell\geq 3$, there exist $k(x)\in\mathbb{N}$, $m(x)\in\mathbb{N}$, $n_{1}(x),n_{2}(x),\ldots,n_{\ell}(x)\in\mathbb{N}$ such that $k(x)+m(x),k(x)+2m(x),\ldots,k(x)+\ell m(x)\in S$ and $a_{n_{1}}(x)=k(x)+m(x),a_{n_{2}}(x)=k(x)+2m(x),\ldots,a_{n_{\ell}}(x)=k(x)+\ell m(x)$. Equations (1.2) ensure that all these integers are taken uniformly on the set $E_{S}$ of Hausdorff dimension $1/2$. Moreover, from the uniformity in multiple recurrence (see [1, Theorem F2]) and Furstenberg’s correspondence principle, the common difference $m(x)$ $(x\in E_{S})$ of the $\ell$-term arithmetic progression can be taken to be bounded from above by a universal constant that depends only on $\ell$ and $\overline{d}(S)$. Even if the upper density of $S$ is zero, similar statements follow from (1.4) and (1.5) provided $S$ has positive upper Banach density.
@@ 1.1.2. Polynomial progressions in $\mathbb{N}$
We proceed to giving more examples of uniform transfer via Theorem 1.1. Bergelson and Leibman established a multiple recurrence theorem with polynomial times [2, Theorem ${\rm A}_{0}$], and as its corollary derived a density combinatorics theorem [2, Theorem ${\rm B}_{0}$] which in particular asserts that any subset of $\mathbb{Z}$ with positive upper Banach density contains ‘arbitrary polynomial progressions’. Their theorem is a vast generalization of Szemerédi’s theorem. For an application to the continued fraction, we need the following version of Bergelson-Leibman’s theorem in $\mathbb{N}$. For completeness we include a proof in Appendix A. Let $\mathbb{Z}_{0}[X]$ denote the set of polynomials with integer coefficients that vanish at $X=0$.
Theorem 1.2 (cf. [2] Theorem ${\rm B}_{0}$).
If $S\subset\mathbb{N}$ and $d_{\rm B}(S)>0$, then for any finite collection $h_{1},h_{2}\ldots,h_{\ell}$ of elements of $\mathbb{Z}_{0}[X]$ there exist $k\in\mathbb{N}$, $m\in\mathbb{N}$ such that $k+h_{1}(m),k+h_{2}(m),\ldots,k+h_{\ell}(m)\in S.$
@@
If $S\subset\mathbb{N}$ and $\overline{d}(S)>0$ (resp. $d_{\rm B}(S)>0$), then for any finite collection $h_{1},h_{2},\ldots,h_{\ell}$ of elements of $\mathbb{Z}_{0}[X]$ there exist [^1] A close inspection into the general construction in §$\S$3.2 shows that in the $\ell$-term arithmetic progression case $h_{1}(X)=X,h_{2}(X)=2X,\ldots,h_{\ell}(X)=\ell X$ one can choose the indices $n_{1},n_{2},\ldots,n_{\ell}\in\mathbb{N}$ so that $n_{1}<n_{2}<\cdots<n_{\ell}$. Similar remarks apply to Theorems 1.6, 1.9 and 1.11. $k\in\mathbb{N}$, $m\in\mathbb{N}$, $n_{1},n_{2}\ldots,n_{\ell}\in\mathbb{N}$ such that $k+h_{1}(m),k+h_{2}(m),\ldots,k+h_{\ell}(m)\in S$ and $a_{n_{1}}(x)=k+h_{1}(m),a_{n_{2}}(x)=k+h_{2}(m),\ldots,a_{n_{\ell}}(x)=k+h_{\ell}(m)$ for any $x\in E_{S}$ (resp. $x\in F_{S}$).
@@ 1.1.3. Second main result
Let $S$ be an infinite subset of $\mathbb{N}$ that has polynomial density with exponent $\alpha$. For any $A\subset S$ with $\overline{d}(A|S)>0$, there exists a subset $E_{S,A}$ of $\{x\in E\colon\{a_{n}(x)\colon n\in\mathbb{N}\}\subset S\}$ such that
@@
dimH{ :xE:xE (1.8)
Since $\mathbb{N}$ has polynomial density with $\alpha=1$, $\beta=0$, Theorem 1.4 includes Theorem 1.1(a) as a special case. Equation (1.8) is a ‘pointwise version’ of the first equality in (1.6), and allows us to immediately transfer density combinatorics statements on subsets of $S$ to statements in dimension theory of continued fractions. This transfer can be done uniformly on $E_{S,A}$ by virtue of (1.7), as developed below.
@@ 1.1.4. Polynomial progressions in primes
The first application of Theorem 1.4 is to the set $\mathbb{P}$ of prime numbers. It is well-known (see e.g., [23, Corollary 3.4]) that the Banach density of $\mathbb{P}$ is zero. Green and Tao [14, Theorem 1.2] proved that any subset of $\mathbb{P}$ with positive upper relative density contains an $\ell$-term arithmetic progression for every $\ell\geq 3$, which they call ‘Szemerédi’s theorem in prime numbers’. Green-Tao’s theorem is a corollary of the wide open conjecture of Erdős and Turán [7] which states that any subset of $\mathbb{N}$ with divergent reciprocal sum contains arithmetic progressions of arbitrary lengths. Subsequently, Tao and Ziegler [36] extended Green-Tao’s theorem to polynomial progressions as follows.
Theorem 1.5 ([36] Theorem 1.3).
If $A\subset\mathbb{P}$ and $\overline{d}(A|\mathbb{P})>0$, then for any finite collection $h_{1},h_{2},\ldots,h_{\ell}$ of elements of $\mathbb{Z}_{0}[X]$ there exist $k\in\mathbb{Z}$, $m\in\mathbb{N}$ such that $k+h_{1}(m),k+h_{2}(m),\ldots,k+h_{\ell}(m)\in A.$
@@
If $A\subset\mathbb{P}$ and $\overline{d}(A|\mathbb{P})>0$, then for any finite collection $h_{1},h_{2},\ldots,h_{\ell}$ of elements of $\mathbb{Z}_{0}[X]$ there exist $k\in\mathbb{Z}$, $m\in\mathbb{N}$, $n_{1},n_{2},\ldots,n_{\ell}\in\mathbb{N}$ such that $k+h_{1}(m),k+h_{2}(m),\ldots,k+h_{\ell}(m)\in A$ and $a_{n_{1}}(x)=k+h_{1}(m),a_{n_{2}}(x)=k+h_{2}(m),\ldots,a_{n_{\ell}}(x)=k+h_{\ell}(m)$ for any $x\in E_{\mathbb{P},A}.$
@@
Surprisingly, there are few literature on continued fractions with prime digits. Schindler and Zweimüller [33] investigated probabilistic laws governing the appearance of prime regular continued fraction digits for irrationals which are typical in the sense of the Lebesgue measure on $(0,1)$. Das and Simmons [4] investigated the set of irrationals in $(0,1)$ all whose regular continued fraction digits are prime numbers.
@@ 1.1.5. Arithmetic progressions in primes of quadratic form
We give three more examples of application of Theorem 1.4, one of which is still related to prime numbers. Let $\mathbb{P}(1)$ denote the set of prime numbers represented by a quadratic form $x^{2}+y^{2}+1$, i.e.,
@@
Linnik [20] proved that $\mathbb{P}(1)$ contains infinitely many primes. Motohashi [25] proved an upper bound #((1)[$\#(\mathbb{P}(1)\cap[1,N])\leq CN/(\log N)^{\frac{3}{2}}$, and Iwaniec [17, Theorem 1] obtained a lower bound of the same order, namely $\mathbb{P}(1)$ has polynomial density with $\alpha=1$, $\beta=3/2$. For more information on $\mathbb{P}(1)$, see [21, 41] for example.
Teräväinen [37] proved that ‘almost every’ even positive integer $n\not\equiv 5,8$ mod $9$ can be written as the sum of two elements of $\mathbb{P}(1)$. As a by-product of a proof of this, he also proved that any subset of $\{p\in\mathbb{P}\colon p=x^{2}+y^{2}+1\ \text{\rm for some coprime }x,y\in\mathbb{N}\}$ with positive relative upper density contains infinitely many non-trivial $3$-term arithmetic progressions. Sun and Pan [34] extended Green-Tao’s theorem to $\mathbb{P}(1)$.
Theorem 1.8 ([34] Theorem 1.1).
@@
If $A\subset\mathbb{P}(1)$ and $\overline{d}(A|\mathbb{P}(1))>0$, then for any integer $\ell\geq 3$ there exist $k\in\mathbb{Z}$, $m\in\mathbb{N}$, $n_{1},n_{2},\ldots,n_{\ell}\in\mathbb{N}$ such that $k+m,k+2m,\ldots,k+\ell m\in A$ and $a_{n_{1}}(x)=k+m,a_{n_{2}}(x)=k+2m,\ldots,a_{n_{\ell}}(x)=k+\ell m$ for any $x\in E_{\mathbb{P}(1),A}.$

### v2

# Density combinatorics theorems in fractal dimension theory of continued fractions

@@ Abstract.
We build a bridge from density combinatorics to fractal dimension theory of continued fractions. We establish a fractal transference principle that transfers common properties of subsets of $\mathbb{N}$ with positive upper density to properties of subsets of irrationals in $(0,1)$ for which the set $\{a_{n}(x)\colon n\in\mathbb{N}\}$ of partial quotients induces an injection $n\in\mathbb{N}\mapsto a_{n}(x)\in\mathbb{N}$. Let ()(*) be a certain property that holds for any subset of $\mathbb{N}$ with positive upper density. The principle asserts that for any $S\subset\mathbb{N}$ with positive upper density, there exists a set $E_{S}\subset(0,1)$ of Hausdorff dimension $1/2$ such that the set $\bigcup_{n\in\mathbb{N}}\bigcap_{x\in E_{S}}\{a_{n}(x)\}\cap S$ has the same upper density as that of $S$, and thus inherits property ()(*). Examples of ()(*) include the existence of arithmetic progressions of arbitrary lengths and the existence of arbitrary polynomial progressions, known as Szemerédi’s and Bergelson-Leibman’s theorems respectively. In the same spirit, we establish a relativized version of the principle applicable to the primes, to the primes of the form $y^{2}+z^{2}+1$ with $y$, $z\in\mathbb{N}$, to the sets given by the Piatetski-Shapiro sequences.
@@ 1. Introduction
$\overline{d}(S)=\limsup_{N\to\infty}\frac{\#({S\cap[1,N]})}{N},$
@@
$d_{\rm B}(S)=\limsup_{N\to\infty}\frac{\sup_{M\in\mathbb{N}}\#({S\cap[M,M+N]})}{N}.$
Note that $d_{\rm B}(S)\geq\overline{d}(S)$. The positive upper density assumption in Szemerédi’s theorem can actually be weakened to the positivity of upper Banach density. Furstenberg [10, 11] established the multiple recurrence theorem and the correspondence principle, linking the field of density combinatorics to recurrence in ergodic theory. Using this principle he obtained a new proof of Szemerédi’s theorem. This development laid a foundation for further extensions of Szemeredi’s theorem to several different directions, see [2, 12, 15, 37, 39, 41] for example. The aim of this paper is to provide counterparts of these density combinatorics theorems in the context of fractal dimension theory of continued fractions.
Each irrational number $x$ in the interval $(0,1)$ has the unique infinite regular continued fraction expansion
$x=\frac{\displaystyle{\hfill{1}\hfill\;\vrule}}{\displaystyle{\vrule\;\hfill{a_{1}(x)}\hfill}}+\frac{\displaystyle{\hfill{1}\hfill\;\vrule}}{\displaystyle{\vrule\;\hfill{a_{2}(x)}\hfill}}+\frac{\displaystyle{\hfill{1}\hfill\;\vrule}}{\displaystyle{\vrule\;\hfill{a_{3}(x)}\hfill}}+\cdots,$ (1.1)
@@
$E=\{x\in(0,1)\setminus\mathbb{Q}\colon a_{m}(x)\neq a_{n}(x)\ \text{for all }m,n\in\mathbb{N}\text{ with }m\neq n\}.$
An investigation of this set was suggested by Erdős in the early stage of fractal dimension theory of continued fractions. Ramharter [35] proved that $E$ is of Hausdorff dimension $1/2$. Ramharter’s result is a refinement of Good’s landmark theorem [14, Theorem 1] which states that the set $\{x\in(0,1)\setminus\mathbb{Q}\colon a_{n}(x)\to\infty\text{ as }n\to\infty\}$ is of Hausdorff dimension $1/2$.
@@ 1.1. Statements of results and their consequences
We now state our two main results (Theorems 1.1 and 1.4), along with consequential theorems (Theorems 1.3, 1.6, 1.9, 1.11 and 1.12) which are manifestations of the density combinatorics theorems [2, 15, 37, 39, 41] in the regular continued fraction.
@@ 1.1.1. First main result
The second equation in (1.2) asserts that the set of elements of $S$ that appear in the sequence of partial quotients of any $x\in E_{S}$ in the same position is of upper density $\overline{d}(S)$. As a ‘pointwise’ version of this, (1.3) asserts that the condition $\overline{d}(\{a_{n}(x)\colon n\in\mathbb{N}\}\cap S)=\overline{d}(S)$ does not cause a dimension drop in $E$.
Equation (1.3) allows us to immediately transfer statements on subsets of $\mathbb{N}$ with positive upper density to statements on continued fractions of points in sets of Hausdorff dimension $1/2$. For example, from (1.3) and Szemerédi’s theorem it follows that if $S\subset\mathbb{N}$ and $\overline{d}(S)>0$, then for any $x\in(0,1)\setminus\mathbb{Q}$ in a set of Hausdorff dimension $1/2$ and for every integer $\ell\geq 3$, there exist $k(x)\in\mathbb{N}$, $m(x)\in\mathbb{N}$, $n_{1}(x),n_{2}(x),\ldots,n_{\ell}(x)\in\mathbb{N}$ such that
$k(x)+m(x),k(x)+2m(x),\ldots,k(x)+\ell m(x)\in S,$
and
$a_{n_{1}(x)}(x)=k(x)+m(x),a_{n_{2}(x)}(x)=k(x)+2m(x),\ldots,a_{n_{\ell}(x)}(x)=k(x)+\ell m(x).$
Equations (1.2) ensure that all the integers $k(x),m(x),n_{1}(x),n_{2}(x),\ldots,n_{\ell}(x)$ are taken uniformly on the set $E_{S}$ of Hausdorff dimension $1/2$. Moreover, a close inspection into Section 3.4 involving the construction of $E_{S}$ shows that one can choose the indices $n_{1}(x),n_{2}(x),\ldots,n_{\ell}(x)$ so that $n_{1}(x)<n_{2}(x)<\cdots<n_{\ell}(x)$. Furthermore, from the uniformity in multiple recurrence (see [1, Theorem F2]) and Furstenberg’s correspondence principle, the common difference $m(x)$ $(x\in E_{S})$ of the $\ell$-term arithmetic progression can be taken to be bounded from above by a universal constant that depends only on $\ell$ and $\overline{d}(S)$. If $\overline{d}(S)=0$ and $d_{\rm B}(S)>0$, then similar statements follow from (1.4) and (1.5). This uniformity of integers and the monotonicity of indices apply to Theorems 1.3, 1.6, 1.9, 1.11 and 1.12.
@@ 1.1.2. Polynomial progressions in $\mathbb{N}$
We proceed to more examples of uniform transfer via Theorem 1.1. Bergelson and Leibman established a multiple recurrence theorem with polynomial times [2, Theorem ${\rm A}_{0}$], and as its corollary derived a density combinatorics theorem [2, Theorem ${\rm B}_{0}$] which in particular asserts that any subset of $\mathbb{Z}$ with positive upper Banach density contains ‘arbitrary polynomial progressions’. Their theorem is a vast generalization of Szemerédi’s theorem. For an application to the continued fraction, we need the following version of Bergelson-Leibman’s theorem in $\mathbb{N}$. For completeness we include a proof in Appendix A. Let $\mathbb{Z}_{0}[X]$ denote the set of polynomials with integer coefficients that vanish at $X=0$.
Theorem 1.2 (cf. [2, Theorem ${\rm B}_{0}$]).
If $S\subset\mathbb{N}$ and $d_{\rm B}(S)>0$, then for any finite collection $h_{1},h_{2}\ldots,h_{\ell}$ of elements of $\mathbb{Z}_{0}[X]$ there exist $k\in\mathbb{N}$, $m\in\mathbb{N}$ such that
$k+h_{1}(m),k+h_{2}(m),\ldots,k+h_{\ell}(m)\in S.$
@@
If $S\subset\mathbb{N}$ and $\overline{d}(S)>0$ (resp. $d_{\rm B}(S)>0$), then for any finite collection $h_{1},h_{2},\ldots,h_{\ell}$ of elements of $\mathbb{Z}_{0}[X]$ there exist $k\in\mathbb{N}$, $m\in\mathbb{N}$, $n_{1},n_{2}\ldots,n_{\ell}\in\mathbb{N}$ such that
$k+h_{1}(m),k+h_{2}(m),\ldots,k+h_{\ell}(m)\in S,$
and for any $x\in E_{S}$ (resp. $x\in F_{S}$),
$a_{n_{1}}(x)=k+h_{1}(m),a_{n_{2}}(x)=k+h_{2}(m),\ldots,a_{n_{\ell}}(x)=k+h_{\ell}(m).$
@@ 1.1.3. Second main result
Let $S$ be an infinite subset of $\mathbb{N}$ that has polynomial density with exponent $\alpha\geq 1$. For any $A\subset S$ with $\overline{d}(A|S)>0$, there exists a subset $E_{S,A}$ of $\{x\in E\colon\{a_{n}(x)\colon n\in\mathbb{N}\}\subset S\}$ such that
@@
dimH{ :xE (1.8)
Equation (1.7) asserts that the set of elements of $A$ that appear in the sequence of partial quotients of any $x\in E_{S,A}$ in the same position is of relative upper density $\overline{d}(A|S)$. As a ‘pointwise’ version of this, (1.8) asserts that the condition $\overline{d}(\{a_{n}(x)\colon n\in\mathbb{N}\}\cap A|S)=\overline{d}(A|S)$ does not cause a dimension drop in $E$.
Since $\mathbb{N}$ has polynomial density with $\alpha=1$, $\beta=0$, Theorem 1.4 includes Theorem 1.1(a) as a special case. Equation (1.8) is a ‘pointwise’ version of the first equality in (1.6), and allows us to immediately transfer density combinatorics statements on subsets of $S$ to statements in fractal dimension theory of continued fractions. This transfer can be done uniformly on $E_{S,A}$ by virtue of (1.7), as developed below.
@@ 1.1.4. Polynomial progressions in primes
The first application of Theorem 1.4 is to the set $\mathbb{P}$ of prime numbers. It is well-known (see e.g., [27, Corollary 3.4]) that the upper Banach density of $\mathbb{P}$ is zero. Green and Tao [15, Theorem 1.2] proved that any subset of $\mathbb{P}$ with positive upper relative density contains an $\ell$-term arithmetic progression for every $\ell\geq 3$, which they call ‘Szemerédi’s theorem in prime numbers’. Green-Tao’s theorem is a corollary of the wide open conjecture of Erdős and Turán [7] which states that any subset of $\mathbb{N}$ with divergent reciprocal sum contains arithmetic progressions of arbitrary lengths. Subsequently, Tao and Ziegler [41] extended Green-Tao’s theorem to polynomial progressions as follows.
Theorem 1.5 ([41, Theorem 1.3]).
If $A\subset\mathbb{P}$ and $\overline{d}(A|\mathbb{P})>0$, then for any finite collection $h_{1},h_{2},\ldots,h_{\ell}$ of elements of $\mathbb{Z}_{0}[X]$ there exist $k\in\mathbb{Z}$, $m\in\mathbb{N}$ such that
$k+h_{1}(m),k+h_{2}(m),\ldots,k+h_{\ell}(m)\in A.$
@@
If $A\subset\mathbb{P}$ and $\overline{d}(A|\mathbb{P})>0$, then for any finite collection $h_{1},h_{2},\ldots,h_{\ell}$ of elements of $\mathbb{Z}_{0}[X]$ there exist $k\in\mathbb{Z}$, $m\in\mathbb{N}$, $n_{1},n_{2},\ldots,n_{\ell}\in\mathbb{N}$ such that
$k+h_{1}(m),k+h_{2}(m),\ldots,k+h_{\ell}(m)\in A,$
and for any $x\in E_{\mathbb{P},A},$
$a_{n_{1}}(x)=k+h_{1}(m),a_{n_{2}}(x)=k+h_{2}(m),\ldots,a_{n_{\ell}}(x)=k+h_{\ell}(m).$
@@
Studying the set of continued fractions with restrictions involving the set of prime numbers is not new, see [4, 13, 25, 38] for example. Mauldin and Urbański [25], Das and Simmons [4] investigated the set of irrationals in $(0,1)$ all whose regular continued fraction digits are prime numbers. Schindler and Zweimüller [38] investigated probabilistic laws governing the appearance of prime regular continued fraction digits for irrationals which are typical in the sense of the Lebesgue measure on $(0,1)$. González Robert et al. [13] computed the Lebesgue measure and the Hausdorff dimension of certain sets of irrationals in $(0,1)$ with infinitely many large prime digits.
@@ 1.1.5. Arithmetic progressions in primes of quadratic form
In this subsection we give three more examples of application of Theorem 1.4, one of which is still related to prime numbers. Let $\mathbb{P}(1)$ denote the set of prime numbers represented by a quadratic form $x^{2}+y^{2}+1$, i.e.,
@@
Linnik [23] proved that $\mathbb{P}(1)$ contains infinitely many primes. Motohashi [29] proved an upper bound $(\mathbb{P}(1)\cap[1,N])\leq CN/(\log N)^{\frac{3}{2}}$, and Iwaniec [18, Theorem 1] obtained a lower bound of the same order, namely $\mathbb{P}(1)$ has polynomial density with $\alpha=1$, $\beta=3/2$. For more information on $\mathbb{P}(1)$, see [24, 46] for example.
Teräväinen [42] proved that ‘almost every’ even positive integer $n\not\equiv 5,8$ mod $9$ can be written as the sum of two elements of $\mathbb{P}(1)$. As a by-product of a proof of this, he also proved that any subset of the set $\{p\in\mathbb{P}\colon p=x^{2}+y^{2}+1\ \text{\rm for some coprime }x,y\in\mathbb{N}\}$ with positive relative upper density contains infinitely many non-trivial $3$-term arithmetic progressions. Sun and Pan [39] extended Green-Tao’s theorem to $\mathbb{P}(1)$.
Theorem 1.8 ([39, Theorem 1.1]).
@@
If $A\subset\mathbb{P}(1)$ and $\overline{d}(A|\mathbb{P}(1))>0$, then for any integer $\ell\geq 3$ there exist $k\in\mathbb{Z}$, $m\in\mathbb{N}$, $n_{1},n_{2},\ldots,n_{\ell}\in\mathbb{N}$ such that
$k+m,k+2m,\ldots,k+\ell m\in A,$
and for any $x\in E_{\mathbb{P}(1),A}$,
$a_{n_{1}}(x)=k+m,a_{n_{2}}(x)=k+2m,\ldots,a_{n_{\ell}}(x)=k+\ell m.$